no. of atoms in 88gm of CO2 ?

Grade:12

11 Answers

Avinash
8 years ago
No. of moles = wt/GMwt =88/44 = 2

No. of atoms or moleules = 2 x Avagadro no. =2 mol x 6.022 x 1023mol-1=12.044 x 1023mol-1
vindhya
8 years ago
thanks a lot. But will the no. of atoms and no.of molecules be same ?
Please let me know, sir.
Avinash
8 years ago
Yes,no. of atoms and no.of molecules be same since it contains avagadro no of molecules or avagadro no of atoms that mutliplied with no. of moles. so the count will be equal
vindhya
8 years ago
thanks sir
Prajwal Kavad
8 years ago
there are 2 moles of CO2 so it has 2times avagadro number of molecules and each molecule has 3 atoms therefore there are total 6xNatoms in 88 grams of CO2
 
plz approve if u r satisfied
Prajwal Kavad
8 years ago
the answer by @avinash is wrong  plz check it sir
vindhya
8 years ago
prajwal kavad, can you please explain me how one molecule has 2 atoms.
i'll be really thankful.
And also tell me how Avinash sir is wrong?
thanks
Avinash
8 years ago
Above ans is for monoatom
Vindhya plz check the correct one

No. of moles = wt/GMwt =88/44 = 2

No. of moleules = 2 x Avagadro no. = 2 mol x 6.022 x 1023mol-1 = 12.044 x 1023


No. of atoms = (1 carbon atom + 2 oxygen atoms) x total no. of molecules
=3 x12.044 x 1023= 36.132 x1023
vindhya
8 years ago
 
thanks a lot to prajwal kavad  for correction and of course Avinash sir  
for a good explanation.
Prajwal Kavad
8 years ago
one molecule has 3 atoms since there is 1C(carbon) atom and 2O(oxygen) atoms
vindhya
8 years ago
ok